OpenStudy (habbababba1):

cost: $38 percent of markup: 58% a. $22.04 c. $15.96 b. $60.04 d. $1.59

OpenStudy (gracygirl):

You are trying to find the new price? Or how much is added on?

OpenStudy (habbababba1):

Find the markup: cost: $38 percent of markup: 58% a. $22.04 c. $15.96 b. $60.04 d. $1.59

OpenStudy (gracygirl):

So what it cost before it was marked up? I'm slightly confused.

OpenStudy (habbababba1):

i am to thats why i posted it im not completely sure

OpenStudy (gracygirl):

Interesting. Let's see... @AloneS @Awolflover1

alones (alones):

Well so Multiply Cost * Mark-up rate \(\ $38 * 0.58\) \[ \frac{ 58 % }{ 100 } =.58\] \(\ $22.04\)

alones (alones):

@habbababba1

OpenStudy (ajspeller):

(New Amounts - Original Amount)/(Original Amount) = Markup or MarkDown

OpenStudy (ajspeller):

\[\frac{ new-original }{ original } = price Change\]